Delectable Parsnip, Inc.s, net income for the most recent year was $10,512. The tax rate was 34 percent. The firm paid $3,587 in total interest expense and deducted $5,673 in depreciation expense What was the companys taxable income for the year? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Taxable income What was the companys EBIT for the year? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) EBIT What was the companys cash coverage ratio for the year? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16) Cash coverage ratio times
Answer #1
Taxable income = Net income / (1-tax%) =10512/(1-34%)
15927.27
EBIT = Taxable income + Interest expense =15927.27+3587
19514.27
Cash Coverage ratio = ( EBIT + Depreciation ) / Interest =(19514.27+5673)/3587
7.02